OpenStudy (anonymous):

Write the equation of the line in standard form that passes through the points (2,-3) and (-1,-4)

OpenStudy (anonymous):

@Nnesha @SanjanaP @zepdrix

Nnesha (nnesha):

so do you know the slope formula

OpenStudy (anonymous):

yes, i found the slope to be 1/3

Nnesha (nnesha):

wowoo good so do you know how to find y-intercept

OpenStudy (anonymous):

i got -16/3?

OpenStudy (anonymous):

i just need help putting it in to standard form

Nnesha (nnesha):

how you get -16/3 seem not good to me

OpenStudy (anonymous):

using the first set of coordinates, y+3=1/3(x-2) and i got y+3=1/3x-7/3 and when i subtracted 3 from doth sides i got -16/3

Nnesha (nnesha):

i'm not understand you add 3 with y you should substitute  suppose to be like this